# Polyelectrolyte adsorption at the solid-liquid interface favors receding contact line instability

### 01 Apr 2026 v1

@@ Abstract
Controlling the motion of non-Newtonian drops on surfaces is crucial for applications ranging from inkjet printing to biomedical devices and food processing. While the macroscopic behavior of viscoelastic drops sliding on tilted hydrophobic surfaces has been characterized—showing reduced velocities and elongation compared to Newtonian fluids—the underlying microscopic mechanisms remain poorly understood. To address this gap, we developed a high-speed, high-resolution reflection microscope that enables direct visualization of the contact line of sliding drops. We used water-soluble polyelectrolyte solutions based on polyacrylamide and let drops sliding on hydrophobic substrates composed of Teflon AF- and PDMS-coated glass slides. The substrate tilting angle was varied between 20° and 45°. We reveal how viscoelasticity influences the dynamics of the receding contact line and drop motion. Our experiments demonstrate that viscoelasticity can destabilize the receding contact line, triggering filament formation. This instability previously observed in the coating of thin viscoelastic films, is reported here for the first time in sliding drops. We further highlight the critical role of polymer charge in this process: while cationic and non-ionic polymers promote filament formation, anionic polymers do not, a difference we attribute to the distinct wetting properties of the solutions. In conclusion, we clarify the interplay between rheology, surface interactions, and drop dynamics.
@@ I Introduction
Viscous fingering or Saffman-Taylor instability Saffman and Taylor (1997) is a well studied instability that arises when a less viscous fluid displaces a more viscous fluid, leading to the formation of ”fingers”. Fingering instabilities also occur at contact lines. For example, at advancing contact lines, when a fluid front is flowing down a tilted plate, regions with larger height flow faster leading to the formation of fingersHuppert (1982). This instability is primarily driven by macroscopic flows, with contact line dynamics playing a minor roleBrenner (1993). More recently, fingering instabilities have also been observed at receding contact lines in the context of non-Newtonian fluids, particularly viscoelastic fluids composed of an aqueous solution of high-molecular weight polymersDeblais et al. (2016). The origin of this instability differs from that at advancing contact lines: it arises from the interplay between Van der Waals and capillary forcesSharma and Wilson (2025). This results in a distinct dependency on contact line velocity. Unlike advancing contact lines, the wavelength of fingers forming at the receding contact line is found to increase with contact line velocity. However, this study does not account for fluid elasticity and does not explain the evolution of filament length with contact line velocity.
@@
A drop sliding across a tilted plate provides a system that allows for the study of both advancing and receding contact lines. The literature on viscoelastic sliding drops is limited and primarily focuses on macroscopic features such as velocity and shape. Varagnolo et al.Varagnolo et al. (2017) showed that sliding viscoelastic drops are elongated compared to Newtonian fluids, with the elongation depending on polymer flexibility and molecular weight. Xu et al.Xu et al. (2016, 2018) demonstrated the deposition of filaments at the rear of drops on superhydrophobic surfaces consisting of pillars. He linked their formation to the interplay between gravity, viscosity, and contact angle hysteresis. However, many aspects remain unclear: How do the physico-chemical properties of sliding drops affect contact line instabilities? Are polymers deposited on the surface during contact line motion? Understanding these questions is valuable for applications that rely on viscoelastic fluids of high molecular weight, including coating processes and transport of biological materials.
@@
In this work, we explore the microscopic deformation of the receding contact lines of viscoelastic drops sliding across a hydrophobic surface. It turned out that viscoelastic sliding drops exhibit receding contact line instabilities leading to the formation of filaments. We demonstrate how the physico-chemical properties of the drop influence the onset of instability and filament formation. We study polyacrylamide (repeat unit -CH2CH(CONH2)-) and two polyacrylamide-based polyelectrolytes. In these, a certain fraction of acrylamide units have been replaced by units bearing a non-zero charge in solution. When in solution, the anionic and cationic polyelectrolytes release their counter-ions and get charged. Polyacrylamide is generally considered to be non-ionic. As shown in Zimmermann et al. (2001); Tandon et al. (2008); Preočanin et al. (2012), zeta-potential measurements indicate that Teflon AF surfaces acquire a negative charge and spontaneoulsy form an electric double layer when exposed to water at neutral pH. We show that the occurrence of instability depends on the charge of the polymer in solution and is due to polymer/surface interactions. Cationic and non-ionic polymers are more prone to develop long filaments while anionic only show small deformation of the receding contact line. Finally, we demonstrate that these filaments lead to the deposition of polymer on the surface.
@@ II.2 Surface preparation
The substrate consisted of glass slides with a thickness of 0.17 mm coated on the backside with a conductive ITO layer (Präzisions Glas & Optik GmbH). Two different hydrophobic surface coatings were pepared. Teflon AF-coated slides were fabricated by dip-coating the glass slides with a 60 nm layer of Teflon-AF from a 1 wt% Teflon AF at a withdrawal speed of 1 cm/min, followed by annealing at 160°C under vacuum for 24 h. PDMS-coated slides were prepared using a grafting-to method.Zhou et al. (2024) The glass slides were drop cast with polydimethylsiloxane (PDMS) with a molecular weight of 6000 g/mol (Sigma Aldrich), placed into an oven at 100°C for 24 h, and subsequently rinsed with toluene and ethanol to remove free PDMS.
@@ III.1 Hydrodynamics of viscoelastic wetting
Regarding the triggering of the instability and filament formation, Sharma et al. Sharma and Wilson (2025) theoretically show that a receding liquid film becomes unstable below a critical thickness. In Newtonian fluids, this leads to the emission of satellite droplets. In viscoelastic fluids, it results in filament formation: the elastic component (i.e., non-zero relaxation time) delays satellite droplet emission and make filament formation possible. However, they did not investigate how the filament characteristics (length, wavelength) depend on the solution. Yet, we observed significant differences in drop motion and contact line morphology depending on the polymer’s charge. We now discuss the origin of these differences.
@@ III.5 Filament dynamics
To understand the role of viscoelasticity in the filament formation, we estimated the magnitude of capillary, viscous and elastic stresses for the case of the cationic polymer. In this case, sliding drops of diameter 4 mm move at typical velocities mm/s. At the bulk scale, the characteristic shear rate of s-1, which matching this shear rate to the rheological curve of the cationic polymer at 2500 ppm concentration (see SI, Fig. S1) shows that the effective viscosity remains roughly four times smaller than the zero-shear value. Pa s.
@@
where and correspond to the normal and radial stresses along the filament and is the extensional rate from Dripping onto surface measurements:
@@
Viscous stresses can be estimated as Pa, which are quite small in comparison with and . Therefore, viscous stresses are negligible as instantaneous tensile resistance strength. Their role is rather dissipative, damping the growth of capillary perturbations. The bead on a tail morphology is thus a result of a Rayleigh-Plateau instability modified by the viscoelasticity, where capillarity drives the process elasticity sustain filament formation and delays breakup, while viscosity governs viscous dissipation. Note that the filaments observed in the “tail on a bead” instability are not stable. Following the argument of Xu et al. Xu et al. (2016, 2018), we compare the breakup time of a viscoelastic filament, , with a convective timescale , where is a characteristic length along the tail. Taking , the experimentally observed bead spacing, and using , we obtain a critical extensional viscosity for the filament survival Pa s. This value is consistent with our observations: the extensional viscosity as determined from DoS is larger than this threshold, and the filaments, although they clearly form, break up rapidly after only a few bead spacings. In contrast to previous studies on micropillared superhydrophobic surfaces, where the bead spacing is constrained by the pillar pitch and filaments are anchored on pinned droplets Xu et al. (2016, 2018), the spacing in our experiments is not imposed by surface topography (PDMS and Teflon AF are smoother). Instead, it reflects the intrinsic wavelength selected by the Rayleigh–Plateau instability, whose growth is delayed by viscoelasticity but not suppressed.
@@ IV Conclusion
Our study demonstrates that viscoelastic sliding drops exhibit receding contact line instabilities, which lead to filament formation. The onset and extent of this instability depend critically on the charge of the polymer: cationic and non-ionic polymers promote the formation of long filaments, whereas anionic polymers show minimal deformation of the contact line. This behavior can be attributed to electrostatic interactions at the solid–liquid interface, leading to polymer adsorption in the case of cationic systems and depletion for anionic ones. Notably, the polymer initially classified as non-ionic was found to be partially protonated and to adsorb onto the surface, behaving similarly to cationic polymers. These findings emphasize the interplay between rheology, surface interactions, and drop dynamics, with potential applications in coating, printing, and surface engineering.
@@ Acknowledgements.
The authors would like to thank Andreas Hanewald and Gunnar Glasser respectively for rheology and SEM measurements, as well as Lin Jian for providing the PDMS-coated samples. We also thank PolyChemie GmbH for providing the polymers used in the study. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 883631) (H.-J. B.).V Author Declarations

### 20 Apr 2026 v2

@@ Abstract
Controlling the motion of non-Newtonian drops on surfaces is crucial for applications ranging from inkjet printing to biomedical devices and food processing. While the macroscopic behavior of viscoelastic drops sliding on tilted hydrophobic surfaces has been characterized—showing reduced velocities and elongation compared to Newtonian fluids—the underlying microscopic mechanisms remain poorly understood. To address this gap, we developed a high-speed, high-resolution reflection microscope that enables direct visualization of the contact line of sliding drops. We used water-soluble polyelectrolyte solutions based on polyacrylamide and let drops sliding on hydrophobic substrates composed of Teflon AF (amorphous fluoropolymer)- and PDMS (Polydimethylsiloxane)-coated glass slides. The substrate tilting angle was varied between 20° and 45°. We reveal how viscoelasticity influences the dynamics of the receding contact line and drop motion. Our experiments demonstrate that viscoelasticity can destabilize the receding contact line, triggering filament formation. This instability previously observed in the coating of thin viscoelastic films, is reported here for sliding drops, for the first time on smooth surfaces. We further highlight the critical role of polymer charge in this process: while cationic and non-ionic polymers promote filament formation, anionic polymers do not, a difference we attribute to the distinct wetting properties of the solutions. In conclusion, we clarify the interplay between rheology, surface interactions, and drop dynamics.
@@ I Introduction
Viscous fingering or Saffman-Taylor instability Saffman and Taylor (1997) is a well studied instability that arises when a less viscous fluid displaces a more viscous fluid, leading to the formation of ”fingers”. Fingering instabilities also occur at contact lines. For example, at advancing contact lines, when a fluid front is flowing down a tilted plate, regions with larger height flow faster leading to the formation of fingersHuppert (1982). This instability is primarily driven by macroscopic fluid dynamics, with contact line dynamics playing a minor roleBrenner (1993). More recently, fingering instabilities have also been observed at receding contact lines in the context of non-Newtonian fluids, particularly viscoelastic fluids composed of an aqueous solution of high-molecular weight polymersDeblais et al. (2016). The origin of this instability differs from that at advancing contact lines: it arises from the interplay between Van der Waals and capillary forcesSharma and Wilson (2025). This results in a distinct dependency on contact line velocity. Unlike advancing contact lines, the wavelength of fingers forming at the receding contact line is found to increase with contact line velocity. However, this study does not account for fluid elasticity and does not explain the evolution of filament length with contact line velocity.
@@
A drop sliding across a tilted plate provides a system that allows for the study of both advancing and receding contact lines. The literature on viscoelastic sliding drops is limited and primarily focuses on macroscopic features such as velocity and shape. Varagnolo et al.Varagnolo et al. (2017) showed that sliding viscoelastic drops are elongated compared to Newtonian fluids, with the elongation depending on polymer flexibility and molecular weight. Xu et al.Xu et al. (2016, 2018) demonstrated the deposition of filaments at the rear of drops on superhydrophobic surfaces consisting of pillars. They linked filament formation to the interplay between gravity, viscosity, and contact angle hysteresis. However, many aspects remain unclear: How do the physico-chemical properties of sliding drops affect contact line instabilities? Are polymers deposited on the surface during contact line motion? Understanding these questions is valuable for applications that rely on viscoelastic fluids of high molecular weight, including coating processes and transport of biological materials.
@@
In this work, we explore the microscopic deformation of the receding contact lines of viscoelastic drops sliding across a hydrophobic surface. It turned out that viscoelastic sliding drops exhibit receding contact line instabilities leading to the formation of filaments. We demonstrate how the physico-chemical properties of the drop influence the onset of instability and filament formation. We study polyacrylamide (repeat unit -CH2CH(CONH2)-) and two polyacrylamide-based polyelectrolytes. In these, a certain fraction of acrylamide units have been replaced by units bearing a non-zero charge in solution. When in solution, the anionic and cationic polyelectrolytes release their counter-ions and get charged. Polyacrylamide is generally considered to be non-ionic. As shown in Zimmermann et al. (2001); Tandon et al. (2008); Preočanin et al. (2012), zeta-potential measurements indicate that Teflon AF (amorphous fluoropolymer) surfaces acquire a negative charge and spontaneoulsy form an electric double layer when exposed to water at neutral pH. We show that the occurrence of instability depends on the charge of the polymer in solution and is due to polymer/surface interactions. Cationic and non-ionic polymers are more prone to develop long filaments while anionic only show small deformation of the receding contact line. Finally, we demonstrate that these filaments lead to the deposition of polymer on the surface.
@@ II.2 Surface preparation
The substrate consisted of glass slides with a thickness of 0.17 mm coated on the backside with a conductive ITO layer (Präzisions Glas & Optik GmbH). Two different hydrophobic surface coatings were pepared. Teflon AF-coated slides were fabricated by dip-coating the glass slides with a 60 nm layer of Teflon AF from a 1 wt% Teflon AF at a withdrawal speed of 1 cm/min, followed by annealing at 160°C under vacuum for 24 h. PDMS (Polydimethylsiloxane)-coated slides were prepared using a grafting-to method.Zhou et al. (2024) The glass slides were drop cast with polydimethylsiloxane (PDMS) with a molecular weight of 6000 g/mol (Sigma Aldrich), placed into an oven at 100°C for 24 h, and subsequently rinsed with toluene and ethanol to remove free PDMS.
@@ III.1 Hydrodynamics of viscoelastic wetting
Regarding the triggering of the instability and filament formation, Sharma et al. Sharma and Wilson (2025) theoretically show that a receding liquid film becomes unstable below a critical thickness due to an interplay between Van der Waals and capillary forces. In Newtonian fluids, this leads to the emission of satellite droplets. In viscoelastic fluids, it results in filament formation: the elastic component (i.e., non-zero relaxation time) delays satellite droplet emission and make filament formation possible. However, they did not investigate how the filament characteristics (length, wavelength) depend on the solution. Yet, we observed significant differences in drop motion and contact line morphology depending on the polymer’s charge. We now discuss the origin of these differences.
@@ III.5 Filament dynamics
To understand the role of viscoelasticity in the filament formation, we estimated the magnitude of capillary, viscous and elastic stresses for the case of the cationic polymer. In this case, sliding drops of diameter 4 mm move at typical velocities mm/s. At the bulk scale, the characteristic shear rate of s-1, which matching this shear rate to the rheological curve of the cationic polymer at 2500 ppm concentration (see SI, Fig. S1) shows that the effective viscosity is quite close to the zero-shear value ( Pa s).
@@
where and correspond to the normal and radial stresses along the filament and is the extensional rate from Dripping onto surface measurements (DoS):
@@
Viscous stresses can be estimated as Pa, which are quite small in comparison with and . Therefore, viscous stresses are negligible as instantaneous tensile resistance strength. Their role is rather dissipative, damping the growth of capillary perturbations. The bead on a tail morphology is thus a result of a Rayleigh-Plateau instability modified by the viscoelasticity, where capillarity drives the process elasticity sustain filament formation and delays breakup, while viscosity governs viscous dissipation. Note that the filaments observed in the “tail on a bead” instability are not stable. Following the argument of Xu et al. Xu et al. (2016, 2018), we compare the breakup time of a viscoelastic filament, , with a convective timescale , where is a characteristic length along the tail. Taking , the experimentally observed bead spacing, and using , we obtain a critical extensional viscosity for the filament survival Pa s. This value is consistent with our observations: the extensional viscosity as determined from DoS is in the same order of magnitude that this threshold, and the filaments, although they clearly form, break up rapidly after only a few bead spacings. In contrast to previous studies on micropillared superhydrophobic surfaces, where the bead spacing is constrained by the pillar pitch and filaments are anchored on pinned droplets Xu et al. (2016, 2018), the spacing in our experiments is not imposed by surface topography (PDMS and Teflon AF are smoother). Instead, it reflects the intrinsic wavelength selected by the Rayleigh–Plateau instability, whose growth is delayed by viscoelasticity but not suppressed.
@@ IV Conclusion
Our study demonstrates that viscoelastic sliding drops exhibit receding contact line instabilities, which lead to filament formation. The onset and extent of this instability depend critically on the charge of the polymer: cationic and non-ionic polymers promote the formation of long filaments, whereas anionic polymers show minimal deformation of the contact line. This behavior can be attributed to electrostatic interactions at the solid–liquid interface, leading to polymer adsorption in the case of cationic systems and depletion for anionic ones, and a modification of the wetting properties of the fluid on the surface. Notably, the polymer initially classified as non-ionic was found to be partially protonated and to adsorb onto the surface, behaving similarly to cationic polymers. These findings emphasize the interplay between rheology, surface interactions, and drop dynamics, with potential applications in coating, printing, and surface engineering.
V Supplementary Material
See the supplementary material for details regarding the solutions’ rheology (flow curves, Fig. S1), SEM image following anionic droplet sliding (Fig. S2), and the sliding of glycerol drops (Fig. S3).
@@ Acknowledgements.
The authors would like to thank Andreas Hanewald and Gunnar Glasser respectively for rheology and SEM measurements, as well as Lin Jian for providing the PDMS-coated samples. We also thank PolyChemie GmbH for providing the polymers used in the study. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 883631) (H.-J. B.) and the Consolidator Grant INTER-ET (grant agreement no. ERC-CoG-2024-101171358) (D. D., O. T.).VI Author Declarations